# VisEscape: A Benchmark for Evaluating Exploration-driven Decision-making in Virtual Escape Rooms

### v2

@@ Abstract
Escape rooms present a unique cognitive challenge that demands exploration-driven planning: players should actively search their environment, continuously update their knowledge based on new discoveries, and connect disparate clues to determine which elements are relevant to their objectives. Motivated by this, we introduce VisEscape, a benchmark of 20 virtual escape rooms specifically designed to evaluate AI models under these challenging conditions, where success depends not only on solving isolated puzzles but also on iteratively constructing and refining spatial-temporal knowledge of a dynamically changing environment. On VisEscape, we observe that even state-of-the-art multimodal models generally fail to escape the rooms, showing considerable variation in their levels of progress and trajectories. To address this issue, we propose VisEscaper, which effectively integrates Memory, Feedback, and ReAct modules, demonstrating significant improvements by performing 3.7 times more effectively and 4.9 times more efficiently on average compared to baseline agents.111[Uncaptioned image] Code: pull-ups/VisEscape
@@ 1 Introduction
Existing works on embodied agents have addressed the importance of exploration, but it has primarily been studied under decomposable and explicit tasks like “Place a cleaned egg in a microwave” [38, 32, 21, 18]. However, the capability of agents to autonomously search for and utilize relevant information from environments where tasks or solutions are not explicitly defined has not been sufficiently investigated.
@@
Figure 2: An illustration of an excerpt of a trajectory from one room in VisEscape. To escape the room successfully, agents must explore multiple directions and diverse views, and interact with various objects. Additionally, they need to infer associations between two or more scenes in different locations to solve creative puzzles.
To bridge this gap, we argue that practical evaluation of decision-making agents should incorporate exploration as a core component. To elicit meaningful exploration, we claim that two key criteria must be met: (1) the state should be partially observable at each timestep, meaning that the agent can only perceive limited aspects of the environment based on its current position and sensors, rather than having full knowledge of the surroundings; and (2) the setup should necessitate a dynamic update of the agent’s state knowledge in response to its actions, ensuring that the effects of those actions—such as modifications to the environment or newly available actions resulting from the acquisition of a new object or removal of constraints—might not be directly observable at the current viewpoint, requiring the agent to actively investigate and continually adapt.
@@
Additionally, VisEscape provides observation as an image for each state and presents executable actions in text form. This brings together the advantages of text-based games, which can define diverse action types (e.g., pouring coffee from a pot or cutting a chain with a nipper), and the benefit of being able to focus on evaluating the agent’s planning or reasoning capabilities of diverse VLMs by isolating the execution of actions from low-level action control.
On VisEscape, we evaluate various VLMs’ performance and observe that even advanced models like Claude-3.5-Sonnet and GPT-4o show low escape rates of only 10%, while all other models achieve 0% without memory management, reflective thinking, or reasoning, prompting us to propose VisEscaper framework, a strong baseline with three modules—Memory, Feedback, and ReAct [45]—enabling effective and efficient problem solving on VisEscape. Equipped with VisEscaper framework, models previously wholly unsuccessful in escaping successfully escaped, while models with pre-existing escape capabilities demonstrate $\times$3.7 enhanced effectiveness and $\times$4.9 improved efficiency. Additionally, we discover that frequent failures of low-performing agents involve repetition of meaningless actions, and that their deficiencies can be mitigated with enhanced reasoning capacity.
@@
2 Related Works
2.1 Interactive Environments for Evaluating LLM/VLM
Traditionally, text-based games have served as a primary benchmark for language models to explore given environments and learn behaviors in an interactive world [8, 17, 33]. More recently, text-based benchmarks have been used as a venue to evaluate various capabilities and tendencies of LLMs in dynamic situations, such as utilizing scientific knowledge [40] and ethical decision-making [31].
Several studies have advanced LLM/VLM-based agents built on Minecraft [39, 25, 11, 16]. In parallel, embodied agent simulators have facilitated evaluating language-driven agents interacting within visually grounded environments [19, 32, 37, 27].
Meanwhile, the GUI environment has recently been proposed as a benchmark [10, 9, 41] for evaluating VLMs’ interaction with visual content. This benchmark was designed to evaluate VLMs’ ability to perceive visual elements on web interfaces[24, 7], reason about their functionality [34], and execute actions to successfully complete given tasks [35, 13, 44, 49].
2.2 LLM/VLM as an Action Decision Maker
Deploying LLM or VLM agents in real-world applications requires effective handling of dynamic and unpredictable environments [29]. Such environments are characterized by changing contexts, unexpected events, and continuous novel observations [30, 47]. To successfully adapt, agents must actively self-evolve using mechanisms for adaptive reasoning and continuous learning [50].
Recent studies have highlighted three critical capabilities for agent adaptation. First, accumulation of experience enables agents to remember previous interactions, particularly failures, and use these experiences to improve future decisions [12]. Second, strategic exploration allows agents to intelligently choose actions based on past outcomes and external feedback, balancing the benefits of learning against potential risks [51]. Third, knowledge abstraction helps agents form generalized high-level representations from raw observations, significantly improving their adaptability in diverse situations [48].
To support these adaptation skills, we integrate three specialized modules into our agent architecture; Memory, Feedback, and ReAct [45]. Together, these modules strengthen agents’ reasoning, planning, and reflection capabilities, enabling effective performance in dynamic settings.
3 VisEscape Construction
3.1 Key Characteristics of VisEscape
@@
1. Self-directed successive task. In all rooms of VisEscape, the only instruction is “escape the room”, without any guidance on how to progress. Under this uncertainty about where to begin, the agent should achieve an average of 8.4 checkpoints (critical tasks that act as bottlenecks in achieving goal completion), which requires an average of 27.2 steps to successfully escape, even for the oracle trajectory. Agents must independently explore, discovering objectives through trial and error. This setup requires them to build and adjust their mental model of the environment, recognizing key objects, spatial relationships, and potential interactions. Success depends on the agent’s ability to hypothesize, test, and adapt their understanding. The task mirrors real-world reasoning challenges, where goals must be inferred, and solutions emerge through engagement, with earlier discoveries becoming relevant later in the process.
2. Diverse interaction with objects. Since all actions available in VisEscape are defined as text, it was possible to define numerous diverse and creative actions that fit with objects placed in the room. For example, an agent can discover a key hidden inside bread by cutting the bread with a knife, open a closet by cutting chains with nippers, and check a hidden code by flipping an hourglass. These interactions are determined by the type of object currently being viewed and the items currently possessed. When an agent performs an action that satisfies the room’s predefined conditions, it triggers state transitions within the room.
Figure 4: An illustration of each component in VisEscape.
@@
3. Associative visual quizzes. One key element of escape rooms is finding connections between different pieces of information, which requires associative thinking—the ability to link seemingly unrelated elements and a fundamental component of creativity [28, 5, 4]. For example, associative thinking helps players connect a symbol on a puzzle with a similar symbol elsewhere and recognize how objects sharing the same color might be related. As escape rooms are ideal for evaluating the associative thinking abilities of VLMs—an area that has remained unexplored—we placed one or two number/letter code locks and corresponding hand-crafted visual quizzes in each room, which require such reasoning. See Figure 3 for examples.
3.2 Dataset Composition
VisEscape consists of 20 distinct rooms. Each room comprises four walls corresponding to four directions, each containing multiple objects that are classified into two types: Receptacles and Items, following the object categorization criteria used in [38]. Receptacles are objects directly located within a room; they can be secured with locks or locking mechanisms and are capable of concealing or containing items. Items can be collected into the inventory, and they serve dual functions: facilitating interaction with other receptacles or providing the necessary clues to unlock the locks. VisEscape has 32 different receptacles and 23 different items, with each room containing a distinctive visual quiz. Each room contains an average of 5.9 receptacles and 5.1 items, along with one or two visual quizzes. Statistics and examples are shown in Table 1 and Figure 4.
3.3 Game Graph
We construct a game graph to model the state of each room in VisEscape providing visual observations corresponding to the current state of the game. The agent can “inspect [object]” to examine receptacles or items more closely by switching to a close-up view, navigate around the room by using “turn to [direction]”, and interact with objects by performing actions defined by each object. Each room has an average of 54.2 unique scenes, which change depending on the actions performed.
3.4 Room Creation Process
To design the room, we utilize three different 3D modeling software programs: Trimble Sketchup, Autodesk Revit, and Chaos Enscape. Specifically, Trimble Sketchup is utilized for object design, Autodesk Revit for object placement and the overall scene composition, and Chaos Enscape for photorealistic rendering. Refer to the Appendix C. for the detailed dataset construction process.
Figure 5: An overview of the three modules comprising VisEscaper, along with examples of inputs and outputs for each module.
4 VisEscaper: Exploring Agent in VisEscape
Figure 5 provides an overview of the VisEscaper framework, which comprises three modules: the Memory, Feedback, and ReAct [45] modules. Each module, along with task-specific instructions, selectively receives information from the environment and responses from other modules, allowing it to efficiently perform tasks by only considering the information essential for that specific module. For the Memory module, since directly using images from multiple turns of observations as input can exceed the model’s context-length capacity, we utilize the same VLM as an image captioner to replace visual observations with textual descriptions.
4.1 Memory module
Since the agent can only observe a limited portion of the room, which causes the entire state of the room to be partially observable, the agent must be equipped with memory capabilities in order to make progress. VisEscaper’s memory consists of three components:
The Memory module of VisEscaper serves a crucial function in mitigating the computational constraints imposed by extended logs during trajectories and addressing the inherent context window constraints of the model while organizing verbose information into structured and prioritized data.
4.2 Feedback module
The Feedback module takes the action performed by the agent along with the preceding and subsequent observations, allowing it to independently analyze the outcomes or impacts of its actions. As illustrated in Figure 5, even if the observations remain unchanged before and after the action, the Feedback module can conclude that “$0$ is not the correct password for this lock.”, which would be difficult to obtain from a single observation alone. Such feedback enables the agent to determine its current actions based on the analysis of the results of its past actions.
4.3 ReAct module
The ReAct module makes final action decisions by integrating the current observation with abstract and structured memory information transmitted from other modules. Inspired by previous research [45], the agent engages in a thinking process prior to action execution, facilitating reasoned decision-making based on the comprehensive information provided. In addition, the agent transmits its reasoning from the previous step to the next step, enabling it to maintain consistent short-term plans across multiple steps.
@@ 5 Experiments
[…]
Claude-3.5-Sonnet VisEscaper 21.7 (x3.44) 32.37 (x2.91) 10.92 (x3.26) 249.5 (x0.87)
BaseAgent 6.3 11.11 3.35 286.7
GPT-4o VisEscaper 13.3 (x4.03) 29.02 (x1.81) 3.91 (x6.51) 276.2 (x0.94)
BaseAgent 3.3 16.00 0.60 294.5
GPT-4o-mini VisEscaper 6.7 21.31 (x3.97) 1.05 291.2 (x0.97)
BaseAgent 0.0 5.37 0.0 300.0
Random 0.0 0.90 0.0 300.0
Table 2: Comparison of random baseline, BaseAgent, and VisEscaper for Claude-3.5, GPT-4o, and GPT-4o-mini. The highest improvement ratios achieved by VisEscaper compared to BaseAgent for each metric are bolded, and the highest scores of the BaseAgent on each metric are underlined.
@@ 5.1 Experiment Setup
For each room, we conducted three different trials and averaged the resulting metrics. To prevent a few challenging puzzles from blocking access to subsequent checkpoints, we experimented with both the default no-hint setting ($exp_{\text{base}}$) and the hint-provided setting ($exp_{\text{hint}}$). In $exp_{\text{hint}}$, if there are no changes in achieved checkpoints for over 30 consecutive turns, a hint guideline message is given to guide the agent toward the next checkpoint. Furthermore, each experiment was terminated early if no significant changes occurred within 100 consecutive turns, and was automatically concluded after a total of 300 steps.
$exp_{\text{base}}$ $exp_{\text{hint}}$
[…]
Claude 3.5 Sonnet 21.7 32.37 10.92 249.5 90.0 96.33 18.87 30.46 103.1
GPT-4o 13.3 29.02 3.91 276.2 95.0 97.39 20.74 26.83 112.9
GPT-4o-mini 6.7 21.31 1.05 291.2 46.7 67.40 32.16 8.93 228.2
InternVL2.5-38B 0.0 6.20 0.0 - 66.7 79.23 41.69 12.23 210.5
InternVL2-40B 0.0 4.91 0.0 - 8.3 48.18 61.61 0.96 291.1
LLaVA-v1.6-34B 0.0 0.54 0.0 - 15.0 34.79 64.97 1.47 289.8
LLaVA-OneVision-7B 0.0 1.25 0.0 - 8.3 41.02 62.89 0.80 294.9
Human 82.5 95.80 51.30 52.8
Random 0.0 0.90 0.00 -
Table 3: The results of diverse VLMs applied to VisEscaper across both $exp_{\text{base}}$ and $exp_{\text{hint}}$. For each experimental setting and metric, the best scores are bolded and the second-best scores are underlined. Experiments without step counts all ended in failure due to reaching the maximum turn limit. For human scores, we assigned four rooms per participant, with a total of 20 participants in the $exp_{\text{base}}$ setting. Refer to Appendix D. for detailed information on the human study.
@@ 5.2 Metrics
We adopted the major metrics widely used in benchmarks for embodied agents.
  • Success Rate (SR): The ratio of escape success in an episode
  • Goal Completion (GC): The ratio of checkpoints completed at the end of an episode
  • Step: Total actions taken by an agent.
  • Success weighted by Path Length [2] (SPL): Success rate weighted by below term:
    Path length of the oracle trajectoryPath length taken by the $\frac{\text{Path length of the oracle trajectory}}{\text{Path length taken by the agent}}$
@@
For $exp_{\text{hint}}$, we define a metric regarding the agent’s hint usage: Hint-assisted Completion Rate (HCR), which denotes the ratio of checkpoints completed by the agent with the assistance of hints. Higher is better for all metrics except for HCR and Step, where lower is better.
5.3 Baseline
For comparison, we conducted controlled experiments with a BaseAgent, which ablates all three modules of VisEscaper on $exp_{\text{base}}$ setting. This experiment evaluated how much progress could be achieved solely with the recent history of actions and observations, excluding the constituent modules. For BaseAgent, only the most recent 20 observations and actions are alternately provided. We omitted open-source models since the success rate of BaseAgent with these models was 0 across all experiments.
According to Table 2, Claude-3.5-Sonnet, GPT-4o, and GPT-4o-mini all demonstrated significant improvements across all metrics-specifically increase $\times$3.7 in success rate, $\times$2.4 in goal completion, $\times$4.9 in SPL, and 7.3% decrease in total steps on average.
Notably, although higher success rates typically require more turns, all models with VisEscaper nevertheless demonstrated a reduction in the average step count, leading to SPL exhibiting the most substantial improvement among all metrics. This result underscores that the implementation of VisEscaper has made the exploration and decision-making processes of the original VLM agent both more effective and efficient.
@@ 5.4 Results
1. Even advanced models cannot achieve success without any hints. Table 3 presents the results of various VLMs integrated within the VisEscaper framework under both $exp_{\text{base}}$ and $exp_{\text{hint}}$. Without any hints ($exp_{\text{base}}$), even Claude-3.5-Sonnet achieves only a 21.7% escape success rate and 32.37% goal completion, which is still 3-4 times lower than the human scores. Considering that humans show a minimal difference between the achievement rate and the final escape success rate (95.80 for GC and 82.5 for SR), the results from all models demonstrate that despite making some progress within each room, they cannot guarantee long-horizontal continuous success when no guidance is provided.
2. High-performing models show a drastic increase in all metrics, achieving objectives with fewer hints and within fewer turns. When hints are provided to stuck agents ($exp_{\text{hint}}$), there is a significant increase in success rate and goal completion across all models. Notably, Claude-3.5-Sonnet and GPT-4o escaped rooms using less than half the average number of steps compared to other models, making them much more efficient at path optimization, which leads to the greatest improvement, especially in SPL, among all metrics. Also, smaller models still improve substantially with hints, but rely on them more frequently, as shown by higher HCR.
3. High-performing models demonstrate accelerated goal completion rates in later stages compared to early stages. Figure 6 illustrates the progression of GC in Table 3 across all rooms during the first 100 steps. Notably, while Claude-3.5-Sonnet performs best overall, it exhibits a distinctly accelerated trajectory after approximately 30-40 steps. This acceleration indicates that strong models better manage accumulated room-state changes, effectively leveraging previous observations to make strategic and efficient decisions later in the trajectory.
Additionally, models exhibiting rapid progress (Claude-3.5-Sonnet and GPT-4o) used fewer hints compared to other models (HCR of $exp_{\text{hint}}$ in Table 3). This indicates that rapid progress by high-performing models does not stem from heavy reliance on hints; instead, they use hints selectively only to resolve specific obstacles and, once past these obstacles, accelerate progress by effectively leveraging information accumulated up to that point.
Figure 6: Average GC (Goal Completion) progression across all rooms and experiments during the first 100 steps. The X-axis represents the number of steps, and the Y-axis represents the goal completion at each step. Solid lines represent results from the $exp_{\text{hint}}$ setting, while dotted lines represent those from the $exp_{\text{base}}$ setting.

### v3

# Instructions for *ACL Proceedings

@@ Abstract
Escape rooms present a unique cognitive challenge that demands exploration-driven planning: players should actively search their environment, continuously update their knowledge based on new discoveries, and connect disparate clues to determine which elements are relevant to their objectives. Motivated by this, we introduce VisEscape, a benchmark of 20 virtual escape rooms specifically designed to evaluate AI models under these challenging conditions, where success depends not only on solving isolated puzzles but also on iteratively constructing and refining spatial-temporal knowledge of a dynamically changing environment. On VisEscape, we observe that even state-of-the-art multimodal models generally fail to escape the rooms, showing considerable variation in their levels of progress and trajectories. To address this issue, we propose VisEscaper, which effectively integrates Memory, Feedback, and ReAct modules, demonstrating significant improvements by performing 3.7 times more effectively and 4.9 times more efficiently on average compared to baseline agents.
Instructions for *ACL Proceedings
Anonymous ACL submission
@@ 1 Introduction
Existing works on embodied agents have addressed the importance of exploration, but it has primarily been studied under decomposable and explicit tasks like “Place a cleaned egg in a microwave” Shridhar_2020_CVPR; puig2018virtualhome; li2021igibson; kim2024realfred. However, the capability of agents to autonomously search for and utilize relevant information from environments where tasks or solutions are not explicitly defined has not been sufficiently investigated.
@@
Figure 2: An illustration of an excerpt of a trajectory from VisEscape. To escape the room successfully, agents must explore multiple directions and diverse views, and interact with various objects. Additionally, they need to infer associations between two or more scenes in different locations to solve creative puzzles.
To bridge this gap, we argue that practical evaluation of decision-making agents should assess their ability to conduct self-directed exploration to identify underlying objectives and subsequently to take reasoned actions. In particular, this entails formulating hypotheses for problem-solving based on information and clues gathered through exploration.
@@
On VisEscape, we evaluate various VLMs’ performance and observe that even advanced models like Claude-3.5-Sonnet and GPT-4o show low escape rates of only 10%, while all other models achieve 0% without memory management, reflective thinking, or reasoning, prompting us to propose VisEscaper framework, a strong baseline with three modules—Memory, Reasoning, and Reflection—enabling effective and efficient problem solving on VisEscape. Equipped with VisEscaper framework, models previously wholly unsuccessful in escaping successfully escaped, while models with pre-existing escape capabilities demonstrate $\times$3.7 enhanced effectiveness and $\times$4.9 improved efficiency. Additionally, [TODO: findings from analysis experiment]
@@
2 Key Challenges of VisEscape
Figure 2 illustrates an example of the problem-solving process required in the VisEscape. The game agent should execute a sequence of tasks comprising diverse interactions.
@@ 2 Key Challenges of VisEscape
1. Self-directed Exploration. As in real escape room games, the game agent in VisEscape receives no instruction except for the instruction ’escape the room’. Therefore, without a clear roadmap, the agent must actively explore and figure out how to proceed through trial and error. This setup requires them to build and adjust their mental model of the environment, recognizing key objects, spatial relationships, and potential interactions. Success depends on the agent’s ability to hypothesize, test, and adapt their understanding. The task mirrors real-world reasoning challenges, where goals must be inferred, and solutions emerge through engagement.
2. Long-Horizontal successive task. Under the uncertainty about how to progress without explicit guidance, the agent should achieve an average of 8.4 checkpoints (critical tasks that act as bottlenecks in achieving goal completion), which requires an average of 27.2 steps to successfully escape, even for the oracle trajectory. The considerable length of the task sequence and the dependencies between checkpoints require agents to effectively leverage information from past trajectory. Success hinges on the ability to integrate these prior discoveries into the ongoing decision-making process.
3 Details on Dataset Composition and Construction
3.1 Dataset Composition
Title. VisEscape consists of 20 distinct rooms. Each room comprises four walls corresponding to four directions, each containing multiple objects that are classified into two types: Receptacles and Items, following the object categorization criteria used in Shridhar_2020_CVPR. Receptacles are objects directly located within a room; they can be secured with locks or locking mechanisms and are capable of concealing or containing items. Items can be collected into the inventory, and they serve dual functions: facilitating interaction with other receptacles or providing the necessary clues to unlock the locks. VisEscape has 32 different receptacles and 23 different items, with each room containing a distinctive visual quiz. Each room contains an average of 5.9 receptacles and 5.1 items, along with one or two visual quizzes. Statistics and examples are shown in Table 1 and Figure 14LABEL:fig:example.
Game Graph. To track the dynamically changing room state and provide corresponding observations, we used a Finite State Machine (FSM). FSM consists of states and transitions-we mapped FSM states to observations matching the current game state and modeled how agent’s actions change observations through FSM state transitions. See Appendix LABEL: for more details.
@@ 3.1 Dataset Composition
Actions. Agent can perform two types of actions: movement and object interaction. For movement, it can “inspect [object]” to examine an object closely or step back to gain a broader view. Additionally, it can use turn to [direction] to view other walls that are not currently observable. After moving, a corresponding observation is given based on the node connections in the game graph. For interaction with objects, the agent can utilize acquired items (e.g., Unlock with key) or perform actions that are set for the objects it can see (§2.2). When an agent performs an action that satisfies the object’s predefined conditions, it triggers state transitions according to game graph. See Appendix LABEL: for examples of diverse object interactions.
Figure 3: Process of VisEscape construction.
3.2 VisEscape Construction
Figure 3 shows the creation process for rooms in VisEscape. (a) Objects are created by 3D modeling software, and those needed for arrangement in each room were sampled. (b) Using the sampled objects, game logic defining the key checkpoints to clear the game was designed with AI-human collaboration. A game graph is then built as a FSM to track the overall game state and enable state transitions triggered by agent’s actions. (c) Objects were arranged in the space according to the room’s setup, and visual observations corresponding to game states were mapped. See Appendix LABEL: for detailed explanations of each process, and the prompts used for game logic design.
Frontier MLLMs Open-source MLLMs
[…]
Human 82.5 95.80 51.30 52.8
Table 2: Performance comparison of various MLLMs and human performance on the VisEscape benchmark. The best results for each metric are bolded. Refer Appendix G for detailed information on the human study. For models that failed to escape entirely, the standard deviation is marked with "-".
@@ 4.1 Experiment Setup
At each step, the model receives the current game state observation, the history of the last 20 observations and actions, held items, and available actions, from which it selects one executable action. For each room, we conducted three different trials and averaged the resulting metrics. Furthermore, each experiment was terminated early if no significant changes occurred within 100 consecutive turns, and was automatically concluded after a total of 300 steps.
@@ 4.2 Metrics
We adopted the major metrics widely used in benchmarks for embodied agents. These include: 1. Success Rate (SR) , which is the ratio of escape success in an episode; 2. Goal Completion (GC) , the ratio of checkpoints completed at the end of an episode; 3. Step, the total actions taken by an agent; and 4. Success weighted by Path Length (SPL) anderson2018evaluation, which is the success rate weighted by the term:
Path length of the oracle trajectoryPath length taken by the $\frac{\text{Path length of the oracle trajectory}}{\text{Path length taken by the agent}}$ (1)
@@
For $exp_{\text{hint}}$, we define a metric regarding the agent’s hint usage: Hint-assisted Completion Rate (HCR), which denotes the ratio of checkpoints completed by the agent with the assistance of hints.
@@ 4.3 Results
Table 2 presents the experimental results for a range of closed-source and open-source VLMs. Without any hints, None of the open-source models achieved a single success. Notably, even sophisticated models such as Claude 3.5 Sonnet and GPT-4o demonstrated success rates below 10 percent. These findings underscore the highly challenging nature of the exploration-centric problem-solving environment provided by VisEscape for current models, concurrently highlighting significant avenues for improvement.
To improve model proficiency in VisEscape, agents must effectively execute two primary capabilities that VisEscape requires: exploration and iterative hypothesis formulation and verification. Rather than devising a specialized framework for the unique ’escape room’ task, we aim to analyze how modules considered crucial in agent modeling contribute to the exploration-centric problem-solving task provided by this benchmark. To this end, we propose VisEscaper, augmented with Memory, Self-feedback, and Reasoning modules, and conduct studies for analysis.
5 Can Memory Management helps Exploration?
In VisEscape, agent can only observe a limited portion of the room-for effective navigation and exploration, the agent should maintain spatial memory that the agent constructs for itself so that it can form a coherent mental map of the space. Additionally, it must remember the actions it has attempted so that it can learn from its trials and avoid redundant efforts.
5.1 Memory module Design
Therefore, we designed and applied a Memory Module to the game agent. The memory module consists of Structured Spatial Memory, Exploration Memory, and Salient Action Memory (Refer to Figure 4), which are constructed and updated periodically. It compressively and structurally manages memory from observation and action histories accumulated over tens to hundreds of steps, while also updating periodically. This allows the agent to make clear decisions from a condensed history, effectively mitigating its processing overload by addressing both cognitive demands and the model’s inherent context window constraints.
5.2 Memory and Effective Exploration
Figure 4: [TODO: how about iteration in figure?]
6 How Game Agents perform reasoning?
The Reasoning module makes final action decisions by integrating the current observation with abstract and structured memory information transmitted from other modules. Inspired by previous research yao2023react, the agent engages in a thinking process prior to action execution, facilitating reasoned decision-making based on the comprehensive information provided. In addition, the agent transmits its reasoning from the previous step to the next step, enabling it to maintain consistent short-term plans across multiple steps.
6.1 Ablation on Reasoning
To analyze how performing reasoning prior to action decisions affects 1) the action decisions themselves and 2) game performance, we conducted an ablation study on the reasoning module.
6.2 Performance of VisEscaper
Model $A_{R}$= $A$ $A_{R}$$\neq$ $A$
$A_{R}$> $A$ $A_{R}$< $A$ ambig.
Claude 3.5 Sonnet 0.47 0.5 0.1 0.4
GPT-4o 0.54 0.1 0.4 0.4
Gemini-1.5-Pro 0.52 0.5 0.1 0.4
GPT-4o-mini 0.53 0.1 0.4 0.4
Table 3: Generated by Spread-LaTeX
7 Memory management helps effective exploration.
7.1 Memory
Memory helps effective exploration.